32.5 ml of H3PO4 requires 15.5 ml of 0.10 M Ca(OH)2. Find M for the H3PO4.
1
Expert's answer
2020-06-15T14:03:05-0400

2H3PO4+3Ca(OH)2=Ca3(PO4)2+6H2O

1.we first find the amount of Ca(OH)2. To do this, we multiply the volume of Ca(OH)2 by the concentration (liter): 0.0155litr*0.1mol/litr=0.00155mol

2.According to the above reaction, we find the amount of H3PO4: 0.00155*2/3=0.001mol

3.now we find the concentration of H3PO4. To do this, we divide the amount of H3PO4 by its volume:

0.001mol/0.0325litr=0.03077≈0.03mol/litr

Answer: 0.03mol/l